It took Terry 7 hours to drive the 390 km from Cold Bay to Hot Lake. For the first part of the trip he averaged 60 km/h, but was forced to slow down for the remainder of the trip to 50km/h.How many hours did Terry spend driving at each speed?
Answer by mananth(16946)

Terry 7 hours to drive the 390 km
first part of the trip he averaged 60 km/h,
the remainder of the trip to 50km/h.
..
let first part distance be x km.
second part distance will be 390-x km
..
first part speed = 60km/hr
second part speed = 50km/hr
..
Time taken in first part + time taken in second part = total time taken
x/60 + 390-x / 50 = 7
LCM = 300
5x+6(390-x) / 300 =7
5x+2340-6x = 2100
-x=-240
x= 240 km
distance =240, speed = 60,
so time = 240/60
= 4 hours in the first part
3 hours in the second part.
..
CHECK
240/60 + 150/5 = 7 hours